# Can light-rings self-gravitate?

### v1

@@ IV Conclusion
In particular, we have considered two different routes to assigning a nonzero mass to the light-ring and that are either based on a discrete concentration of photons on a specific radial location or on suitable distribution of photons in a given region. Notwithstanding a number of differences between the two scenarios, we have found that, in both cases, the inclusion of the energy contribution from the accumulated photons leads to the generation of new light-rings. Such new light-rings can either appear at well-defined but discrete locations, or be fused with zero separation in a well-defined region. Interestingly, and somewhat disappointingly, all light-ring configurations considered here are unstable and a small perturbation, either via the inclusion on an additional photon onto the light-ring or via the absorption of a photon by the black hole, leads to a catastrophic destruction of the light-ring structures and to the dispersion of the photons to spatial infinity or inside the black hole. Hence, the answer to the question in the title is: “yes, light-rings can self-gravitate, but they are dynamically unstable”.
This work could be extended in a number of different ways. First, it would be useful to perform a rigorous perturbative analysis that would provide, among other things, the typical timescale of the instability. Second, although these spacetimes are unstable, the fact that they can be described analytically makes them useful tools to study the response of non-trivial black-hole configurations to various perturbations and explore, for instance, the properties of quasi-normal modes or the emergence of echos as different light-rings are involved in the perturbation. Third, while we do not expect that the outcome would change in the presence of rotation, the work presented here could be extended by considering Kerr spacetimes and the corresponding light-rings there. Finally, the exploration of self-gravity started here does not require the presence of an horizon and can be extended to horizonless ultracompact objects as long as they possess an unstable light-ring. Interestingly, these compact objects also possess stable light-rings Mazur and Mottola (2001); Cunha et al. (2017); Jampolski and Rezzolla (2024); Di Filippo (2024). While the analysis for the unstable light-ring would not deviate from the one studied here, novel and potentially intriguing results could emerge regarding self-gravitating stable light-rings. As a final remark, we note that while the scenarios explored in this paper are so idealised to require the intervention of an advanced civilisation, they actually provide important information on the fundamental properties of realistic light-rings in general relativity.
Acknowledgements
@@
It is a pleasure to thank R. Carballo-Rubio, P. Cunha, C. Ecker, C. Herdeiro, P. Kocherlakota, R. Kumar Walia, D. Jampolski, S. Liberati, and P. Schupp for interesting discussions and suggestions. FDF acknowledges financial support from the PRIMUS/23/SCI/005 and UNCE24/SCI/016 grants by Charles University, and the GAR 23-07457S grant from the Czech Science Foundation. LR acknowledges financial support from the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006), from the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (Grant No. 884631), and from the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair.

### v2

@@ Abstract
In particular, we have considered two different routes for assigning a nonzero mass to the light-ring and that are either based on a discrete concentration of photons on a specific radial location or on suitable distribution of photons in a given region. Notwithstanding a number of differences between the two scenarios, we have found that, in both cases, the inclusion of the energy contribution from the accumulated photons leads to the generation of new light-rings. Such new light-rings can either appear at well-defined but discrete locations, or be fused with zero separation in a well-defined region. Interestingly, and somewhat disappointingly, all light-ring configurations considered here are unstable and a small perturbation, either via the inclusion on an additional photon onto the light-ring or via the absorption of a photon by the black hole, leads to a catastrophic destruction of the light-ring structures and to the dispersion of the photons to spatial infinity or inside the black hole. Hence, the answer to the question in the title is: “yes, light-rings can self-gravitate, but they are dynamically unstable”.
This work could be extended in several ways. First, by performing a perturbative analysis that would also provide the timescale of the instability. Second, by using these analytical spacetimes to study the quasi-normal modes or the emergence of echos. Third, by considering Kerr spacetimes and the corresponding light-rings. Furthermore, much of the work started can be extended to horizonless ultracompact objects as long as they possess an unstable light-ring a hence also stable light-rings Mazur and Mottola (2001); Cunha et al. (2017b); Guo and Gao (2021); Jampolski and Rezzolla (2024); Di Filippo (2024). Finally, while the scenarios explored in this paper are so idealised to require the intervention of an advanced civilisation, future investigations might consider more realistic sources of photons. Even without fine-tuning the impact parameter, photons with an impact parameter close enough to populate the light-ring will orbit the black hole several times, spending on average more time in this region. This would naturally lead to an over-density at the light-ring similar to the one considered in our work.
Acknowledgements.
@@
It is a pleasure to thank R. Carballo-Rubio, P. Cunha, C. Ecker, C. Herdeiro, P. Kocherlakota, R. Kumar Walia, D. Jampolski, S. Liberati, H. Olivares, and P. Schupp for interesting discussions and suggestions. FDF acknowledges financial support from the PRIMUS/23/SCI/005 and UNCE24/SCI/016 grants by Charles University, and the GAR 23-07457S grant from the Czech Science Foundation. LR acknowledges financial support from the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006), from the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (Grant No. 884631), and from the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair.